OpenStudy (anonymous):

Find the slope and y intercept? 5x-2y=8?

hartnn (hartnn):

using y=mx+c where m is slope and c is y intercept....we get slope=2.5 and y-int=4

OpenStudy (anonymous):

show work! @hartnn

OpenStudy (anonymous):

ok, so rearranging: 2y=5x-8 y=(5/2)x-4 y=mx+b where m is slope and b is y intercept so slope is 5/2 and y int = -4

OpenStudy (anonymous):

okay thanks